Let P(x) = x + ax4 + bx3 + cx? + dx +e. P(4) = P(5)=P(6)=P(7)=P(8) = 0. What is the value of a - b+c-d+e? Numerically the ansexplain
Answer #1

For a polynomial equation of degree n ,

  P(x)=x^{n}+a_{1}x^{n-1}+a_{2}x^{n-2}+.................+a_{n}=o , whose roots are \alpha _{1},\alpha _{2}....\alpha _{n}

we have the relation between the roots and the coefficients as :

  • the sum of roots =  -a_{1}
  • the sum of products of roots taken two at a time =  a_{2}
  • the sum of products of roots taken three at a time =  -a_{3}
  • ..............................
  • the sum of products of roots taken n at a time = (-1)^{n}a_{n}

Here we have

P(x)=x^{5}+ax^{4}+bx^{3}+cx^{2}+dx+e where 4 , 5 , 6 , 7 , 8 are the roots of P(x)=0

Thus from the above mentioned relation , we have

  -a= 4+5+6+7+8

\Rightarrow a= (-4)+(-5)+(-6)+(-7)+(-8)

b=4*5+4*6+4*7+4*8+5*6+5*7+5*8+6*7+6*8+7*8

-c=4*5*6+4*5*7+4*5*8+4*6*7+4*6*8+4*7*8+5*6*7+5*6*8+5*7*8+6*7*8

  

d=4*5*6*7+4*5*6*8+4*6*7*8+4*5*7*8+5*6*7*8

  -e=4*5*6*7*8

  

\Rightarrow a-b+c-d+e=(-4)+(-5)+(-6)+(-7)+(-8)+(-4*5)+(-4*6)+(-4*7)+(-4*8)+(-5*6)+(-5*7)+(-5*8)+(-6*7)+(-6*8)+(-7*8)+(-4*5*6)+(-4*5*7)+(-4*5*8)+(-4*6*7)+(-4*6*8)+(-4*7*8)+(-5*6*7)+(-5*6*8)+(-5*7*8)+(-6*7*8)+(-4*5*6*7)+(-4*5*6*8)+(-4*6*7*8)+(-4*5*7*8)+(-5*6*7*8)+(-4*5*6*7*8)